OpenStudy (anonymous):

Four years ago, Samantha was three times as old as Danny. Four years from now, Samantha will be only twice as old as Danny. How old are they now ?

OpenStudy (imstuck):

I have come up with this system of equations, and I THINK it is correct. If Danny, 4 years ago, was 3 times younger than Samantha, the first equation would be this: 3(D-4) = S - 4 In four years is Danny is only twice as young as Samantha, the second equation would be this: 2(D+4) = S + 4 If you solve the first one for S and then sub in that expression for S in the second equation, you get this equation: 2D + 8 = 3D - 8 + 4 -D = -12 so Danny is 12 and Samantha is 28.

OpenStudy (anonymous):

simplify \[[(\frac{ 5}{ 6 }^{3})]\times \left[ \left( \frac{ 5 }{ 6 } \right) \right]\]

OpenStudy (anonymous):

add three